Circular No. 6537

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997D IN NGC 1536 Duilia de Mello, Observatorio Nacional, Rio de Janeiro, reports her

discovery of a supernova on Jan. 14.15 UT with the European Southern Observatory (ESO) 1.52-m telescope at La Silla; SN 1997D is located 11′′ east and 43′′ south of the nucleus of NGC 1536 (α = 4^{h}11^{m}.0, δ = −56^{o}29′, equinox 2000.0). S. Benetti, ESO, writes: “Inspection of a fully-reduced spectrum (range 369.5–757.8 nm, resolution ∼ 0.6 nm), obtained on Jan. 14.15 with the ESO 1.52-m telescope (+ Boller & Chivens spectrograph) by the discoverer, confirms SN 1997D to be a peculiar type-II supernova, ∼ 100 days after maximum. The spectrum contains a variety of lines superimposed on a red continuum. Hα shows an unusual double-P-Cyg profile, from which expansion velocities of ∼ 5000 and 1200 km/s are derived (the redshift of the galaxy being 1296 km/s). A line with P-Cyg profile (peak measured at 663.3 nm and minimum at 661.5 nm) is visible on the Hα red wing. The Na I-D doublet is present with a P-Cyg profile (expansion velocity of ∼ 3700 km/s), together with a strong P-Cyg line of Ba II 614.2-nm (expansion velocity ∼ 2200 km/s) and P-Cyg lines of Sc II at 552.7, 565.8, and 624.5 nm (expansion velocity of ∼ 1800 km/s). Emission lines of Fe II at 492.4, 501.8, and 516.9 nm are present in the supernova spectrum. A relatively strong emission line at 734.3 nm has been identified with the Ca II] 729.1–732.3-nm doublet. Hβ is very weak if present at all.” An R image obtained by G. Massone (Osservatorio di Torino) with the 0.9-m Dutch telescope on Jan. 15.05 yields a supernova magnitude of 16.3.

SUPERNOVA 1997B IN IC 438 M. Cavagna, Sormano, Italy, reports the following precise position of

SN 1997B from observations by P. Sicoli and V. Giuliani with the 0.5-m reflector + CCD on Jan. 14.844 UT: α = 5^{h}53^{m}01^{s}.95, δ = −17^{o}51′55′′.7 (equinox 2000.0).

SUPERNOVA 1997C IN NGC 3160 W.-d. Li, Beijing Astronomical Observatory (BAO), reports that the

time of the spectrogram reported on IAUC 6536 was Jan. 14.7 UT. Additional prediscovery observations of SN 1997C have been identified in the BAO data, with the following unfiltered CCD magnitudes available: 1996 Dec. 23 UT, 17.8; 27, 17.3.

1997 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green